# Cross-Service Threat Intelligence in LLM Services using Privacy-Preserving Fingerprints

### v1

@@ Abstract
We present BinaryShield, the first privacy-preserving threat intelligence system that enables secure sharing of attack fingerprints across compliance boundaries. BinaryShield transforms suspicious prompts through a unique pipeline combining PII redaction, semantic embedding, binary quantization, and randomized response mechanism to potentially generate non-invertible fingerprints that preserve attack patterns while providing privacy. Our evaluations demonstrate that BinaryShield achieves an F1-score of 0.94, significantly outperforming SimHash (0.77), the privacy-preserving baseline, while achieving 64x storage reduction and 38x faster similarity search compared to dense embeddings.
1 Introduction
@@
The rapid adoption of Large Language Models (LLMs) in digital services is driving a paradigm shift in human-computer interaction. Companies now operate many LLM-based services across diverse domains, collectively handling billions of queries each day. Typically, organizations like Microsoft maintain multiple, logically isolated (siloed) LLM services, such as enterprise AI assistants, consumer chat applications, API based LLM services, and developer-focused coding agents. These services are siloed because they are tailored to meet distinct business objectives and end-user requirements, necessitating separate operational boundaries. Each service has distinct model stacks, logging pipelines, and compliance boundaries with strong internal data governance policies to protect user data and privacy. In addition to internal privacy regulations, these services are also subject to governmental regulations such as GDPR and HIPAA, which mandate strict data handling practices to protect user privacy and sensitive information. While this siloed architecture is crucial for maintaining operational independence, it has an unintended consequence: fragmented security telemetry and a weakened collective defense. With little correlation or data sharing between services, organizations are left with a disjointed security posture as each service operates independently, responding to threats in isolation. This fragmentation creates a significant challenge for incident response teams, as they lack a unified view of the threat landscape across their LLM services.
@@
Motivation and Problem Statement. The current defense paradigm is compartmentalized, with each service responding independently to detected attacks. When a new attack vector is identified (typically through media reports, user complaints, or post-breach analysis) security teams patch their systems (e.g., update prompts classifiers). However, this knowledge remains confined to the affected service. Furthermore, attacks are inevitable because existing defenses are probabilistic and can be bypassed (Debenedetti et al., 2025; Costa et al., 2025). Thus, organizations need retrospective systems to pinpoint the perpetrator and timing of incidents.
[…]
Our Contributions. We propose the concept of threat intelligence across compliance boundaries in LLM services. We materialize this idea in BinaryShield, the first privacy-preserving fingerprinting technique for prompt injection attacks that enables secure information sharing between otherwise siloed services. When a potential threat (prompt injection) is detected within any service by its defense mechanism, BinaryShield generates a non-reversible, lightweight fingerprint that distills the semantic essence of the malicious prompt, making it possible to share actionable threat signals while preserving privacy.
Our approach balances the need for threat intelligence sharing with user privacy and regulatory compliance. BinaryShield achieves this through a multi-stage process that progressively removes identifying information while retaining the essential semantic characteristics for effective threat detection. The process begins with personally identifiable information (PII) redaction, where BinaryShield removes sensitive data such as social security numbers, names, and other identifying markers from input prompts. This initial privacy layer is crucial for protecting users whose prompts may be incorrectly flagged as malicious by automated defenses. However, PII redaction alone is insufficient, as the remaining text may still contain contextual information that could compromise user privacy when shared externally. To capture the semantic essence of potentially malicious prompts, BinaryShield generates high-dimensional embeddings from the redacted text. These embeddings, which represent text as dense floating-point vectors, achieve state-of-the-art performance in natural language processing tasks by encoding semantic relationships in vector space. However, embeddings are not secure for sharing across compliance boundaries. Recent research demonstrates that embeddings pose significant privacy risks, as they can allow adversaries to reconstruct original input text (Walsh, ; Li et al., 2023; Tragoudaras et al., 2025; OWASP2025, 2025). Moreover, traditional embeddings present practical challenges for large-scale threat intelligence system. Each dimension typically requires 32 bits of storage, and similarity computations using cosine similarity or dot products demand substantial computational resources, often necessitating GPU acceleration for efficient threat matching.
BinaryShield addresses these fundamental challenges through two innovative non-reversible transformations by exploiting insights from quantization and differential privacy. First, BinaryShield performs binary quantization on the floating-point embeddings, converting each dimension to a single bit based on its sign, assigning 1 for positive values and 0 for negative values. This quantization achieves remarkable efficiency gains, reducing storage requirements while simultaneously enhancing privacy. The key insight underlying this approach is that by discarding magnitude information and retaining only directional information, we make prompt reconstruction exponentially more difficult while preserving sufficient semantic structure for threat sharing and detection. To render original prompt reconstruction practically impossible, BinaryShield applies differential privacy to the binary embeddings. Leveraging the insights from the principle of randomized response (Warner, 1965; Erlingsson et al., 2014; Wang et al., 2017), BinaryShield independently flips each bit of the binary vector with a precisely calibrated probability. This noise addition fundamentally alters the binary vector representation while preserving its calibrated utility for threat correlation, ensuring that adversaries cannot reverse-engineer the original prompt from the shared fingerprint. The resulting binary vector is efficient in storage and computation, enabling rapid Hamming distance similarity checks. More importantly, after these privacy-preserving transformations, the resulting binary vector becomes safe to share across compliance boundaries, enabling effective cross-service threat intelligence without compromising user privacy or regulatory compliance.
Evaluations. We conduct comprehensive evaluation of BinaryShield spanning threat correlation effectiveness, privacy calibration, scalability, and operational efficiency (Section 3). BinaryShield significantly outperforms privacy-preserving baseline across attack variants (17-point F1 advantage on sophisticated attacks) and exhibits smooth privacy-utility trade-offs with theoretical noise alignment. In real-world enterprise-scale scenarios, BinaryShield achieves 79.2% threat correlation accuracy (93% of non-private baseline performance) alongside efficiency gains (38x search speedup, 64x storage reduction).
2 BinaryShield Design
The proliferation of LLM-based services within organizations creates a fundamental tension between security effectiveness and privacy compliance. While traditional malware defense systems successfully share threat signatures across boundaries, no comparable mechanism exists for natural language-based attacks on LLMs. BinaryShield addresses this gap by enabling organizations to share actionable threat intelligence about prompt injection attacks without violating privacy regulations or exposing sensitive user data.
The system is designed to achieve four primary objectives: (1) Semantic Preservation: capture the essential characteristics of malicious prompts that enable cross-service threat correlation; (2) Privacy Guarantee: ensure computational non-invertibility such that original prompts cannot be reconstructed from shared fingerprints hence upholding compliance with internal and government regulations (i.e. GDPR); (3) Computational Efficiency: enable efficient fingerprint generation and search; and (4) Operational Compliance: maintain strict separation of user data across compliance boundaries while enabling collaborative defense. To this end, BinaryShield transforms potentially malicious prompts through a series of privacy-preserving operations, ultimately producing compact binary fingerprints that can be safely shared between services while maintaining regulatory compliance. The key insight underlying our approach is that prompt injection attacks, despite surface-level variations, share semantic patterns that can be captured and compared without revealing the original content. By combining techniques from data anonymization, natural language processing, quantization and differential privacy, we create non-reversible fingerprints that preserve these essential semantic characteristics while providing privacy guarantees against reconstruction attacks.
[…]
@@ 2 BinaryShield Design
Figure 2: BinaryShield approach for privacy-preserving fingerprint generation for cross-service threat intelligence sharing. The pipeline progressively transforms potential prompt injections through PII redaction, semantic embedding, binary quantization, and differential privacy to produce shareable fingerprints that preserve privacy while enabling threat correlation across compliance boundaries.
High-Level Architecture.
Figure 1 illustrates the complete BinaryShield system architecture. An attacker issues semantically related prompts ($x,y,z$) to the services of an enterprise organization that reside in separate compliance boundaries (Figure 1-➊). Each service applies their own defense mechanism ($D1$, $D2$, and $D3$) to incoming prompts. Service-2 flags a prompt-injection attempt (Figure 1-➋). After this step, BinaryShield’s fingerprinting pipeline is invoked in the service’s compliance boundary (Figure 1-➌). At this stage, the attack log is separated into two components. The first and the most critical component is private information: this contains the sensitive content of the prompt itself, holding the essential details of the attack. Although critical for accurate event correlation, this data must remain within the service’s compliance boundary to safeguard privacy due to regulatory requirements. The suspicious prompt then undergoes a series of privacy-preserving transformations to generate a non-reversible privacy fingerprint (hash) within the attacked service’s compliance boundary (Figure 1-➌). A detailed description of the privacy fingerprint generation process appears in Section 2.1. The second part consists of non-private system metadata, such as the tools used during response generation and the geographical region. Since this metadata contains no sensitive data, it can be safely shared among peer services as part of complete signature. Sharing these features enhances attack correlation by providing complementary evidence to surface relevant patterns pointing to attack source and channel, which helps reduce false positives and prioritize alerts. For instance, matching system metadata across services, such as both accessing an email tool, significantly increases the confidence in the correlation of potential attacks. Concatenating both parts forms a composite attack fingerprint that discloses no readable text (Figure 1-➍) This encoded hash is then securely broadcast to peer services for threat correlation (Figure 1-➎). Note that peer services (i.e., Services 1 and 3) use the exact same technique to fingerprint their local private queries and system metadata as used in Step ➌ of Figure 1. For instance, after searching their logs, Service 1 and Service 3 independently find two and one similar attacks, respectively (Figure 1-➏). Policy-driven response actions are then executed for each service based on the correlation results (Figure 1-➐). Note that there is currently no public dataset available for system metadata. Therefore, our evaluation of BinaryShield focuses exclusively on private information (prompt) fingerprinting. Nonetheless, including system metadata features in the overall design demonstrates the full architectural intent of BinaryShield and highlights additional layers of security and context.
2.1 BinaryShield One-Way Fingerprint Generation Design
@@
Output: Privacy-preserving fingerprint $f$ broadcast to $\mathcal{S}$; correlated alerts at peers
@@ 15  Append $\tilde{b}_{i}$ to $\tilde{\mathbf{b}}$ ;
  // Send fingerprint to all peers
18
Algorithm 1 BinaryShield: Privacy-Preserving Threat Intelligence Pipeline
We consider an adversary with several capabilities: possesses white-box access, meaning they have complete knowledge of the BinaryShield algorithm, including the embedding models, quantization methods, and privacy parameters. The adversary’s primary objective is to reconstruct original prompts from shared fingerprint, potentially exposing sensitive user information. Our system must, therefore, prevent such reconstruction while preserving utility for effective threat detection. Based on our threat model and regulatory constraints, BinaryShield must satisfy several critical requirements. First, it must ensure privacy preservation by making fingerprints computationally non-invertible, thereby preventing the reconstruction of original prompts even in the case of white-box access. Second, semantic preservation remains essential; despite privacy transformations, the fingerprints must retain enough semantic information to enable the identification of similar attack patterns. Third, the system must maintain computational efficiency, capable of processing millions of queries each day with minimal latency impact. Fourth, it must offer cross-boundary compatibility, allowing fingerprints to be shared across services with varying compliance requirements without violating relevant regulations. Finally, the system must support approximate matching, enabling similarity searches that can detect related attacks despite minor variations in the input. Figure 2 and Algorithm 1 depicts the BinaryShield pipeline for generating privacy-preserving fingerprints (hashes) of prompt injection attacks. Below, we detail each component of the pipeline, explaining how they collectively satisfy our design requirements.
2.1.1 PII Redaction
@@
The initial stage of the BinaryShield pipeline directly targets privacy risks by identifying and removing personally identifiable information (PII) that may be present in flagged prompts. While automated defense mechanisms aim to identify malicious content, they may inadvertently flag benign prompts containing sensitive user data.
@@
Figure 3: Distribution of PII entities in prompts flagged by BinaryShield. The most common entity include person names.
@@
A key challenge in PII redaction is removing sensitive information while preserving the semantic intent and structure of potential attacks. Our approach addresses this by detecting and redacting structured data such as social security numbers, credit card numbers, phone numbers, email addresses, as well as person names, organizations, locations, and other identifying entities. Identified PII is replaced with generic placeholders that maintain command structure, entity relationships, and linguistic patterns indicative of injection attempts. For example, "Transfer \$5000 from John Smith’s account 123456789" becomes "Transfer [AMOUNT] from [PERSON]’s account [ACCOUNT]". This balance ensures the redacted text retains sufficient information for downstream processing while protecting user privacy from the earliest stage of the pipeline. By design, redaction is intended to be one-way and non-reversible: no original PII should survive beyond this step, ensuring compliance with privacy regulations before any downstream processing. While the effectiveness of redaction depends on the accuracy of underlying detection methods, our approach minimizes the risk of residual PII to the greatest extent possible. BinaryShield uses Microsoft’s Presidio open-source library (Microsoft, ) for PII detection and redaction. Figure 3 reports the PII entities and their respective counts that BinaryShield flags in the prompt injection dataset from (Shen et al., 2024), used in BinaryShield’s evaluation (Section 3). However, even after PII redaction, the remaining text may still contain sensitive or revealing information that cannot be safely shared across compliance boundaries. To address this, BinaryShield’s pipeline generates privacy-preserving semantic fingerprints, enabling robust threat detection without exposing any underlying content.
2.1.2 Semantic Feature Extraction
Following PII redaction, BinaryShield must capture the semantic essence of the potentially malicious prompt in a format suitable for comparison and analysis. To this end, BinaryShield employs state-of-the-art transformer-based language models to generate high-dimensional embeddings that encode the semantic relationships within the text. This transformation serves as the foundation for our semantic preservation requirement. This capture the meaning and intent of the potentially malicious prompt. The embedding process transforms the redacted text into a dense vector $\mathbf{e}\in\mathbb{R}^{d}$, where $d$ is the embedding dimension. The embedding space exhibits several properties that are crucial for threat detection. First, similar attack patterns tend to cluster together in the embedding space, regardless of surface-level variations. Second, the embeddings are robust to paraphrasing, so minor rewording or synonym substitution results in nearby embeddings, enabling the detection of attack variants.
Binary Quantization.
While semantic embeddings effectively capture prompt meaning, they pose significant challenges for privacy-preserving threat intelligence. Dense embeddings can leak information about the original text (Walsh, ; Li et al., 2023; Tragoudaras et al., 2025; OWASP2025, 2025). They also require substantial computational resources for large-scale similarity searches. Our binary quantization addresses both concerns. We transform the continuous embedding $\mathbf{e}\in\mathbb{R}^{d}$ into a binary vector $\mathbf{b}\in\{0,1\}^{d}$ using sign-based quantization:
$b_{i}=\begin{cases}1&\text{if }e_{i}>0\\ 0&\text{otherwise}\end{cases}$
where $e_{i}$ is the $i$-th dimension of the embedding vector and $b_{i}$ is the corresponding bit in the binary vector. This transformation discards magnitude information while preserving the directional (sign) information of each embedding dimension. Binary quantization provides several critical benefits: Irreversibility: The transformation from $\mathbb{R}$ to $\{0,1\}$ is many-to-one, making exact reconstruction impossible. Each bit only indicates whether the original value was positive or negative, discarding all magnitude information. Dimensionality preservation: Despite the dramatic reduction in information per dimension, the high dimensionality ($d$) of the binary embedding preserves the essential geometric properties needed for similarity detection. The sign of each embedding dimension collectively maintains sufficient information to distinguish between semantically similar and dissimilar prompts. Storage efficiency: Each dimension requires only 1 bit instead of 32 bits (float32), achieving a 32x reduction in storage requirements and even more in case of float64. Our empirical evaluation shows that binary quantization preserves semantic relationships and achieves comparable threat detection performance, while providing substantial privacy benefits (Section 3).
@@ 2.1.2 Semantic Feature Extraction
Figure 4: Impact of BinaryShield’s privacy parameter ($\alpha$) on binary vector quality. As $\alpha$ increases, privatized vector (red triangles) move closer to original vector (blue circles), showing improved utility while maintaining differential privacy through controlled bit-flipping (i.e., 1p1-p) noise.
2.1.3 BinaryShield’s Randomized-Response Layer
While data anonymization (PII redaction) and binary quantization significantly enhances privacy, adversaries with auxiliary information might still attempt reconstruction attacks. To close this gap, the pipeline enforces local differential privacy (LDP) on every fingerprint by applying classical randomized response (Warner, 1965; Erlingsson et al., 2014; Wang et al., 2017) independently to each bit before the fingerprint crosses the compliance boundary. Local differential privacy is particularly attractive in cross-boundary threat-intelligence settings because the guarantee is enforced before data leave the originator’s compliance perimeter; no trusted aggregator or secure enclave is required. Even an adversary with white-box knowledge of the entire pipeline and unlimited access to all shared fingerprints cannot violate the worst-case privacy bound. Moreover, in contrast to the global-DP setting, the magnitude of noise required for LDP is independent of the dataset size, allowing BinaryShield to scale to millions of daily fingerprints without ever decreasing privacy protection. Consider a $d$-bit fingerprint $\mathbf{b}=(b_{1},\dots,b_{d})\!\in\!\{0,1\}^{d}$ and let $\alpha\!>\!0$ denote the per-bit privacy budget. Classical randomized response publishes a perturbed vector $\tilde{\mathbf{b}}$ obtained by flipping each bit independently with probability 1p1-p, where
$p=\frac{e^{\alpha}}{e^{\alpha}+1}$

### v2

# arXiv:2509.05608v2[cs.CR] 28 Feb 2026

@@ Abstract
We present BinaryShield, the first privacy-preserving threat intelligence system that enables secure sharing of attack fingerprints across compliance boundaries. BinaryShield transforms suspicious prompts through a unique pipeline combining PII redaction, semantic embedding, binary quantization, and randomized response mechanism to potentially generate privacy-preserving fingerprints that preserve attack patterns while providing privacy. Our evaluations demonstrate that BinaryShield achieves an F1-score of 0.94, significantly outperforming SimHash (0.77), the privacy-preserving baseline, while achieving storage reduction and 38x faster similarity search compared to dense embeddings.
I Introduction
@@
The rapid adoption of Large Language Models (LLMs) in digital services is driving a paradigm shift in human-computer interaction. Companies now operate many LLM-based services across diverse domains, collectively handling billions of queries each day. Typically, organizations like Microsoft maintain multiple, logically isolated (siloed) LLM services, such as enterprise AI assistants, consumer chat applications, API-based LLM services, and developer-focused coding agents. These services are siloed because they are tailored to meet distinct business objectives and end-user requirements, necessitating separate operational boundaries. Each service has distinct model stacks, logging pipelines, and compliance boundaries with strong internal data governance policies to protect user data and privacy. In addition to internal privacy regulations, these services are also subject to governmental regulations such as GDPR and HIPAA, which mandate strict data handling practices to protect user privacy and sensitive information. While this siloed architecture is crucial for maintaining operational independence, it has an unintended consequence: fragmented security telemetry and a weakened collective defense. With little correlation or data sharing between services, organizations are left with a disjointed security posture as each service operates independently, responding to threats in isolation. This fragmentation creates a significant challenge for incident response teams, as they lack a unified view of the threat landscape across their LLM services.
Prompt Injection and Current Defenses. The new interaction model, based on natural language prompts, introduces a critical security risk: prompt injection attacks. Often described as the “SQL injection” of AI, prompt injection is recognized as the top threat in the OWASP 2025 Top 10 for LLMs. Prompt injection exploits LLMs’ inability to distinguish legitimate user instructions from malicious commands embedded in input. These attacks can manipulate system prompts to leak private data, execute malicious code via tool interfaces, or generate harmful content. With emerging technologies like Model Context Protocol (MCP) and autonomous agents, the impact of prompt injection extends beyond text manipulation to arbitrary code execution and potential system compromise [41]. Organizations and researchers have developed various defense mechanisms against these attacks [1, 2]. Furthermore, more than one defense is deployed to enhance security within a compliance boundary [2]. However, a critical limitation stems directly from privacy regulations that prevent the sharing of user prompts between compliance boundaries (i.e., services).
Motivation and Problem Statement. The current defense paradigm is compartmentalized, with each service responding independently to detected attacks. When a new attack vector is identified (typically through media reports, user complaints, or post-breach analysis), security teams patch their systems (e.g., update prompts classifiers). However, this knowledge remains confined to the affected service. Furthermore, attacks are inevitable because existing defenses are probabilistic and can be bypassed [13, 12]. Thus, organizations need retrospective systems to pinpoint the perpetrator and timing of incidents.
This reactive, siloed approach prevents organization-wide visibility into the threat landscape. An attack discovered in one service today may have existed undetected in another for months. Security analysts lack tools to search for similar historical attack patterns across service boundaries, creating a “correlation gap”. As a result, they cannot fully assess the scale and sophistication of an attack, identify its variants, or proactively protect all services from emerging threats. Traditional malware defense addressed similar coordination challenges decades ago by exchanging signatures: antivirus engines share hash or pattern-based fingerprints of malicious binaries without revealing proprietary information [31]. To our knowledge, no comparable, privacy-preserving, and practically deployable threat intelligence mechanism exists for natural-language prompts in LLM services. This paper addresses the following problem: How can an organization securely share threat intelligence about prompt injection attacks across its compliance boundaries without violating privacy regulations?
Ideal Characteristics. A privacy-preserving fingerprinting mechanism for prompt injection attacks is essential for secure information sharing across compliance boundaries. Such a fingerprint must encode the semantic core of the malicious prompt, capturing its essential characteristics beyond surface tokens. It should be privacy preserving, so the original input (prompt) cannot be recovered. The technique must support secure approximate matching for “find-similar” searches across data stores and remain lightweight for real-time generation and distribution. When an attack is detected in one service, the fingerprint can be broadcast to peer services. These peers can then search historical logs for related incidents, flag live traffic, and train local defenses, enabling proactive, collaborative security. Cross-boundary threat intelligence will enable faster patching without compromising user privacy.
Our Contributions. We propose the concept of threat intelligence across compliance boundaries in LLM services. We materialize this idea in BinaryShield, the first privacy-preserving fingerprinting technique for prompt injection attacks that enables secure information sharing between otherwise siloed services. When a service’s defense mechanism detects a potential prompt-injection threat, BinaryShield generates a privacy-preserving, lightweight fingerprint that captures the malicious prompt’s semantic essence, allowing actionable threat signals to be shared without compromising privacy. Note, we use prompt injection as a broader term, similar to its use in the OWASP Top 10 for LLMs, where jailbreaking is treated as a form of prompt injection [35].
Our approach balances the need for threat intelligence sharing with user privacy and regulatory compliance. BinaryShield achieves this through a multi-stage process that progressively removes identifying information while retaining the essential semantic characteristics for effective threat detection. The process begins with personally identifiable information (PII) redaction, where BinaryShield removes sensitive data such as social security numbers, names, and other identifying markers from input prompts. This initial privacy layer is crucial for protecting users whose prompts may be incorrectly flagged as malicious by automated defenses. However, PII redaction alone is insufficient, as the remaining text may still contain contextual information that could compromise user privacy when shared externally. To capture the semantic essence of potentially malicious prompts, BinaryShield generates high-dimensional embeddings from the redacted text. These embeddings, which represent text as dense floating-point vectors, achieve state-of-the-art performance in natural language processing tasks by encoding semantic relationships in vector space. However, embeddings are not secure for sharing across compliance boundaries. Recent research demonstrates that embeddings pose significant privacy risks, as they can allow adversaries to reconstruct original input text [48, 30, 46, 42]. Moreover, traditional embeddings present practical challenges for large-scale threat intelligence system. Each dimension typically requires 32 bits of storage, and similarity computations using cosine similarity or dot products demand substantial computational resources, often necessitating GPU acceleration for efficient threat matching.
BinaryShield addresses these fundamental challenges through two innovative non-reversible transformations by exploiting insights from quantization and differential privacy. First, BinaryShield performs binary quantization on the floating-point embeddings, converting each dimension to a single bit based on its sign, assigning 1 for positive values and 0 for negative values. This quantization achieves remarkable efficiency gains, reducing storage requirements while simultaneously enhancing privacy. The key insight underlying this approach is that by discarding magnitude information and retaining only directional information, we make prompt reconstruction exponentially more difficult while preserving sufficient semantic structure for threat sharing and detection. To render original prompt reconstruction practically impossible, BinaryShield applies differential privacy to the binary embeddings. Leveraging the insights from the principle of randomized response [51, 19, 49], BinaryShield independently flips each bit of the binary vector with a precisely calibrated probability. This noise addition fundamentally alters the binary vector representation while preserving its calibrated utility for threat correlation, ensuring that adversaries cannot reverse-engineer the original prompt from the shared fingerprint. The resulting binary vector is efficient in storage and computation, enabling rapid Hamming distance [37] similarity checks. More importantly, after these privacy-preserving transformations, the resulting binary vector becomes safe to share across compliance boundaries, enabling effective cross-service threat intelligence without compromising user privacy or regulatory compliance.
Evaluations. We conduct comprehensive evaluation of BinaryShield spanning threat correlation effectiveness, privacy calibration, scalability, and operational efficiency (Section III). BinaryShield significantly outperforms privacy-preserving baseline across attack variants (17-point F1 advantage on sophisticated attacks) and exhibits smooth privacy-utility trade-offs with theoretical noise alignment. In real-world enterprise-scale scenarios, BinaryShield achieves 79.2% threat correlation accuracy (93% of non-private baseline performance) alongside efficiency gains (38x search speedup).
II BinaryShield Design
@@
Figure 2: BinaryShield approach for privacy-preserving fingerprint generation for cross-service threat intelligence sharing. The pipeline transforms potential prompt injections through PII redaction, semantic embedding, binary quantization, and differential privacy to produce shareable fingerprints that preserve privacy while enabling threat correlation across compliance boundaries.
@@
Output: Privacy-preserving fingerprint $f$ broadcast to $\mathcal{S}$
@@ 15 Append $\tilde{b}_{i}$ to $\tilde{\mathbf{b}}$ ;
// Send fingerprint to peers
Algorithm 1 BinaryShield
The proliferation of LLM-based services within organizations creates a fundamental tension between security effectiveness and privacy compliance. While traditional malware defense systems successfully share threat signatures across boundaries [31], no comparable mechanism exists for LLM services. BinaryShield addresses this gap by enabling organizations to share actionable threat intelligence about prompt injection attacks without violating privacy regulations.
The system is designed to achieve four primary objectives: (1) Semantic Preservation: capture the essential characteristics of malicious prompts that enable cross-service threat correlation; (2) Privacy Guarantee: ensure that original prompts cannot be reconstructed from shared fingerprints, hence upholding regulations; (3) Computational Efficiency: enable efficient fingerprint generation and search; and (4) Operational Compliance: maintain strict separation of user data across compliance boundaries while enabling collaborative defense. To this end, BinaryShield transforms potentially malicious prompts through a series of privacy-preserving operations, ultimately producing compact binary fingerprints that can be safely shared between services. The key insight underlying our approach is that prompt injection attacks, despite surface-level variations, share semantic patterns that can be captured and compared without revealing the original content.
High-Level Architecture. As a motivating example, consider a company that offers both Enterprise AI and Consumer AI services operating under separate compliance boundaries. When Enterprise AI detects a prompt injection attack, it cannot warn Consumer AI due to regulations prohibiting raw prompt sharing, leaving Consumer AI vulnerable to the same attack. Figure 1 illustrates how BinaryShield addresses this challenge. An attacker issues semantically related prompts ($x,y,z$) to the services of an enterprise organization that reside in separate compliance boundaries (Figure 1-➊). Each service applies their own defense mechanism ($D1$, $D2$, and $D3$) to incoming prompts. Service-2 flags a prompt-injection attempt (Figure 1-➋). After this step, BinaryShield’s fingerprinting pipeline is invoked in the service’s compliance boundary (Figure 1-➌). At this stage, the attack log is separated into two components. The first and most critical component is private information: this contains the sensitive content of the prompt itself, holding the essential details of the attack. Although critical for accurate event correlation, this data must remain within the service’s compliance boundary to safeguard privacy. The suspicious prompt then undergoes a series of privacy-preserving transformations to generate a fingerprint within the attacked service’s compliance boundary (Figure 1-➌). The fingerprint generation process is detailed in Section II-A. The second part consists of non-private system metadata, such as the tools used during response generation and the geographical region. Since this metadata contains no sensitive data, it can be safely shared among peer services as part of complete signature. Sharing these features enhances attack correlation by providing complementary evidence to surface relevant patterns pointing to attack source and channel, which helps reduce false positives and prioritize alerts. For instance, matching system metadata across services, such as both accessing an email tool, significantly increases the confidence in the correlation of potential attacks. Concatenating both parts forms a composite attack fingerprint that discloses no readable text (Figure 1-➍) This fingerprint is then securely broadcast to peer services for threat correlation (Figure 1-➎). Note that peer services (i.e., Services 1 and 3) use the exact same technique to fingerprint their local private queries and system metadata as used in Step ➌ of Figure 1. For instance, after searching their logs, Service 1 and Service 3 independently find two and one similar attacks, respectively (Figure 1-➏). Policy-driven response actions are then executed for each service based on the correlation results (Figure 1-➐). Note that there is currently no public dataset available for system metadata. Therefore, our evaluation of BinaryShield focuses exclusively on private information (prompt) fingerprinting. Nonetheless, including system metadata features in the overall design demonstrates the full architectural intent of BinaryShield and highlights additional layers of security and context.
II-A BinaryShield Fingerprint Generation Design
BinaryShield design is guided by several key requirements. A primary goal is to ensure data privacy by transforming original prompts into fingerprints through a one-way process, making the reconstruction of the original prompt computationally difficult. Simultaneously, these fingerprints must preserve enough semantic information to be useful for identifying related attack patterns via approximate matching. Operationally, the system is built for computational efficiency to handle millions of daily queries with low latency and is designed with cross-boundary compatibility in mind, allowing fingerprints to be shared across services with varying internal regulatory requirements. Algorithm 1 and Figure 2 depicts the BinaryShield’s pipeline for generating privacy-preserving fingerprints of prompt injection attacks. Below, we detail each component of the pipeline.
II-A1 PII Redaction
@@
BinaryShield’s pipeline initial stage directly targets privacy risks by identifying and removing personally identifiable information (PII) that may be present in flagged prompts. While automated defenses aim to identify malicious content, they may inadvertently flag benign prompts containing sensitive user data.
@@
Figure 3: Distribution of PII entities in prompts flagged by BinaryShield. Most common: person names.
@@
The main challenge is removing PII while preserving the semantic structure of potential attacks. Our approach addresses this by detecting and redacting structured data (using [39]) such as social security numbers, credit card numbers, phone numbers, email addresses, person names, organizations, locations, and other identifying entities. Identified PII is replaced with generic placeholders that maintain command structure, entity relationships, and linguistic patterns indicative of injection attempts. For example, “Transfer \$5000 from John Smith’s account 123456789” becomes “Transfer [AMOUNT] from [PERSON]’s account [ACCOUNT]”. This balance ensures the redacted text retains sufficient information for downstream processing while protecting user privacy from the earliest stage of the pipeline. This one-way redaction minimizes PII exposure before any downstream analysis. Figure 3 reports the PII entities and their respective counts that BinaryShield flags in the prompt injection dataset from [45], used in BinaryShield’s evaluation (Section III). However, because redaction alone is insufficient to protect all sensitive information, the pipeline’s next stage generates a privacy-preserving semantic fingerprint.
@@
Figure 4: Impact of BinaryShield’s privacy parameter ($\alpha$). As $\alpha$ increases, privatized vector (red triangles) move closer to original vector, showing improved utility while maintaining differential privacy through controlled bit-flipping (i.e., 1p1-p) noise.
II-A2 Semantic Feature Extraction
Following PII redaction, BinaryShield must capture the semantic essence of the potentially malicious prompt in a format suitable for comparison and analysis. To this end, BinaryShield employs state-of-the-art embedding models [50, 23] to generate high-dimensional embeddings that encode the semantic relationships within the text. This transformation serves as the foundation for our semantic preservation requirement. This captures the meaning and intent of the potentially malicious prompt. The embedding process transforms the redacted text into a dense vector $\mathbf{e}\in\mathbb{R}^{d}$, where $d$ is the embedding dimension. The embedding space exhibits several properties that are crucial for threat detection. First, similar attack patterns tend to cluster together in the embedding space, regardless of surface-level variations. Second, the embeddings are robust to paraphrasing, so minor rewording or synonym substitution results in nearby embeddings, enabling the detection of attack variants.
Binary Quantization. While semantic embeddings effectively capture prompt meaning, they pose significant challenges for privacy-preserving threat intelligence. Dense embeddings can leak information about the original text [48, 30, 46, 42]. They also require substantial computational resources for large-scale similarity searches. Our binary quantization addresses both concerns. We transform the continuous embedding $\mathbf{e}\in\mathbb{R}^{d}$ into a binary vector $\mathbf{b}\in\{0,1\}^{d}$ using sign-based quantization:
$b_{i}=\begin{cases}1&\text{if }e_{i}>0\\ 0&\text{otherwise}\end{cases},$
where $e_{i}$ is the $i$-th dimension of the embedding vector and $b_{i}$ is the corresponding bit in the binary vector. This transformation provides several critical benefits by design. First, it acts as a lossy compression mechanism. By mapping each continuous value in $\mathbb{R}$ to a binary state in $\{0,1\}$ based on its sign, the process discards all magnitude information. This many-to-one mapping makes it computationally difficult to infer the original floating-point vectors, thus enhancing data protection. Second, despite this information reduction, the high dimensionality ($d$) of the binary embedding preserves essential geometric properties needed for similarity detection. Finally, this method is highly storage-efficient, reducing the memory footprint for each dimension from 32 bits (for a float32) to just 1 bit, a 32x reduction.
II-A3 BinaryShield’s Randomized-Response Layer
To build upon the baseline protections of data anonymization and binary quantization, our pipeline incorporates Local Differential Privacy (LDP) for a more formal privacy guarantee. By applying a randomized response mechanism [51, 19, 49] to each bit before a fingerprint crosses the compliance boundary, we gain several critical advantages. This approach provides a strong, mathematically-defined upper bound on privacy loss for each fingerprint, ensuring plausible deniability with a formal guarantee of privacy regardless of any auxiliary information a person might possess. The guarantee is enforced locally before data leaves its origin, which is ideal for cross-boundary threat intelligence as it removes the need for a trusted central aggregator. Furthermore, unlike in the global DP, the magnitude of the required noise is independent of the dataset size, ensuring our privacy protection remains constant and does not degrade as the system scales to millions of daily fingerprints.
Consider a $d$-bit fingerprint $\mathbf{b}=(b_{1},\dots,b_{d})\!\in\!\{0,1\}^{d}$ and let $\alpha\!>\!0$ denote the per-bit privacy budget. Classical randomized response publishes a perturbed vector $\tilde{\mathbf{b}}$ obtained by flipping each bit independently with probability 1p1-p, where
$p=\frac{e^{\alpha}}{e^{\alpha}+1}.$